# From Complex-Analytic Models to Sparse Domination: A Dyadic Approach of Hypersingular Operators via Bourgain’s Interpolation Method

### v2

@@ 3. $\bullet$
Acknowledgement. The authors are grateful to Árpád Bényi for drawing their attention to strongly singular Calderón–Zygmund operators. They also thank Yongsheng Han, Cody Stockdale and Yun–Hao Lee for helpful discussions during various stages of this work. The first author was supported by the Simons Travel grant MPS-TSM-00007213.
@@ 4. Critical line estimates for the hypersingular Bergman projection
where in the last estimate, we used the standard integral estimate [29, Theorem 1.12]. Therefore, for any $\alpha>0$, as in Lemma 3.4,
@@
which gives the desired weak type bounds. ∎
Our next goal is to prove estimates at the other endpoint $\big(\frac{1}{3-2t},1\big)$. To begin with, we have the following unboundedness result.
@@
On the other hand, by [29, Theorem 1.12], one has
@@
First, recall that for any $z,w\in{\mathbb{D}}$, there exists a Carleson tent $Q_{J}$, containing both $z$ and $w$ such that $|Q_{J}|\simeq\left|1-z\overline{w}\right|^{2}$ (see, e.g., [24, Section 3]). Then using the $1/3$-trick, one can find some $I\in{\mathcal{D}}$ or $\widetilde{{\mathcal{D}}}$ (see, Remark 1.4), such that
@@ 5. Hypersingular sparse operators
    $\{E(Q)\}_{Q\in{\mathcal{S}}}$ are mutually disjoint.
We begin with a prototypical model case. Let $t>1$, and for simplicity, assume that $\mathcal{S}$ be an $\eta$-sparse family of cubes in $\mathbb{R}^{n}$ such that $Q\subseteq Q_{0}=[0,1]^{n}$ for all $Q\in{\mathcal{S}}$. We define the (pointwise) hypersingular sparse operator associated with $\mathcal{S}$ by
@@ 5.1. Graded family
$\bullet$ First, let ${\mathcal{G}}^{\{0\}}$ denote the collection of all maximal dyadic cubes in $\mathcal{G}$}. For simplicity, we may assume that ${\mathcal{G}}^{\{0\}}=\{[0,1]^{n}\}$. Otherwise, we decompose $\mathcal{G}$} into finitely many such collections and treat each one separately (by translation and dilation). To this end, we define $\mathfrak{G}_{0}:=\ell(Q_{0})$, where $\ell(Q)$ denotes the sidelength of a dyadic cube $Q$.
$\bullet$ Next, let ${\mathcal{G}}^{\{1\}}$ be the subcollection of all maximal dyadic cubes in ${\mathcal{G}}\setminus{\mathcal{G}}^{\{0\}}$, and define $\mathfrak{G}_{1}:=\inf_{Q\in{\mathcal{G}}^{\{1\}}}\ell(Q)$. Iterating this procedure, we obtain a decomposition of $\mathcal{G}$} into layers $\{{\mathcal{G}}^{\{j\}}\}_{j\geq 0}$ together with the associated scales $\{\mathfrak{G}_{j}\}_{j\geq 0}$. Observe that for each $j\geq 1$, the dyadic cubes in ${\mathcal{G}}^{\{j\}}$ are mutually disjoint.
@@
Let ${\mathcal{G}}\subseteq{\mathcal{D}}$ be a collection of dyadic cubes in $\mathbb{R}^{n}$ such that ${\mathcal{G}}^{\{0\}}=\{[0,1]^{n}\}$, and write ${\mathcal{G}}=\bigcup_{j\geq 0}{\mathcal{G}}^{\{j\}}$ as above. We say that $\mathcal{G}$} is graded if
@@
Let $\eta\in(0,1)$, ${\mathcal{S}}\subseteq{\mathcal{D}}$ be a graded $\eta$-sparse family in $0,1]^{n}$ with degree $K_{{\mathcal{S}}}$ as in Definition 5.4, $1<t<1-\frac{\log_{2}(1-\eta)}{nK_{{\mathcal{S}}}}$, and $\mathbb{A}_{{\mathcal{S}}}^{t}$ be the associated hypersingular sparse operator. Then the following statements hold.
@@
Observe that, by the sparseness assumption of $\mathcal{S}$}, for any $j\geq 0$, one has
(5.4) $\sum_{Q\in{\mathcal{S}}^{\{j\}}}|Q|\lesssim(1-\eta)^{j},$
@@
(1). First, consider ${\mathcal{S}}^{(0)}=\{[0,1]^{n}\}$. By the sparseness of $\mathcal{S}$},
$\mathbb{A}_{{\mathcal{S}}}^{t}f(z)\leq\frac{1}{\left|[0,1]^{n}\right|^{t}}\int_{[0,1]^{n}}|f(x)|dx\leq 1,\qquad z\in E([0,1]^{n}),$
where $E([0,1]^{n})$ is defined as in Definition 5.4. Since $|E([0,1]^{n})|>\eta$, $\mathbb{A}_{{\mathcal{S}}}^{t}f(z)$ can only take larger values on $0,1]^{n}\backslash E([0,1]^{n})$, whose size is at most $1-\eta$.
(2). Next, we consider the next layer ${\mathcal{S}}^{(1)}$. Using the sparseness of $\mathcal{S}$} again, we find that
$\mathbb{A}_{{\mathcal{S}}}^{t}f(z)\leq 1+\frac{1}{|Q|^{t}}\int_{Q}|f(x)|dx\leq 1+|Q|^{1-t}\leq 1+2^{nK_{{\mathcal{S}}}(t-1)},\quad z\in\bigcup_{Q\in{\mathcal{S}}_{1}}E(Q)$
By the fact that the cubes in ${\mathcal{S}}^{(1)}$ are mutually disjoint and (5.4), we have
$\left|\bigcup_{Q\in{\mathcal{S}}^{(1)}}E(Q)\right|=\sum_{Q\in{\mathcal{S}}^{(1)}}|E(Q)|\leq\sum_{Q\in{\mathcal{S}}^{(1)}}|Q|\lesssim(1-\eta)^{j}.$
Moreover, $\mathbb{A}_{{\mathcal{S}}}^{t}f(z)$ can only take large values on
$0,1]^{n}\backslash\left(E([0,1]^{n})\cup\bigcup_{Q\in{\mathcal{S}}^{(1)}}E(Q)\right),$
whose size is at most $(1-\eta)^{2}$.
@@
where $C_{1},C_{2}>0$ are some absolute constants [^6] Here, we can take $C_{1}=\frac{2^{nK_{{\mathcal{S}}}(t-1)}}{2^{nK_{{\mathcal{S}}}(t-1)}-1}$ and $C_{2}=\frac{1}{2^{nK_{{\mathcal{S}}}(t-1)}-1}$. that only depend on $n,K_{{\mathcal{S}}}$ and $t$, then $z\in\bigcup_{\ell=0}^{J}\bigcup_{Q\in{\mathcal{S}}^{(\ell)}}E(Q)$, whose size is at least $1-(1-\eta)^{J+1}$. Therefore, for any $\alpha>0$ sufficiently large, if $\mathbb{A}_{{\mathcal{S}}}^{t}f(z)>C_{1}\alpha-C_{2}$, then
$z\in[0,1]\backslash\bigcup_{\ell=0}^{\widetilde{J}}\bigcup_{Q\in{\mathcal{S}}^{(\ell)}}E(Q)$
@@ 6. Some open problems
where ${\mathbb{B}}_{n}$ is the unit ball in $\mathbb{C}^{n}$ and $dV$ is the normalized volume measure on ${\mathbb{B}}_{n}$. In [27], Zhao and Zhou characterized the strong-type $L^{p}_{\alpha}({\mathbb{B}}_{n})\to L^{q}_{\beta}({\mathbb{B}}_{n})$ bounds for these operators under various assumptions on the parameters $a,b,c,\alpha$, and $\beta$. Here, for $1\leq p<\infty$ and $1<\alpha<\infty$, the space $L^{p}_{\alpha}({\mathbb{B}}_{n})$ is defined with respect to the measure $dv_{\alpha}(z)=c_{\alpha}(1-|z|^{2})^{\alpha}\,dV(z)$, where $c_{\alpha}$ is chosen so that $v_{\alpha}({\mathbb{B}}_{n})=1$. A natural question is whether the method developed in the present paper can be applied to obtain critical-line estimates for Forelli–Rudin type operators in the hypersingular regime (see, e.g., [27, Theorem 1.1]).

### v3

@@ 3. $\bullet$
Acknowledgement. The authors are grateful to Árpád Bényi for drawing their attention to strongly singular Calderón–Zygmund operators. They also thank Yongsheng Han, Cody Stockdale, Yun–Hao Lee, Kenan Zhang, and Zipeng Wang for helpful discussions during various stages of this work. The first author was supported by the Simons Travel grant MPS-TSM-00007213.
@@ 4. Critical line estimates for the hypersingular Bergman projection
where in the last estimate, we used the standard integral estimate [31, Theorem 1.12]. Therefore, for any $\alpha>0$, as in Lemma 3.4,
@@
which yields the desired weak-type bounds. ∎
Our next goal is to establish estimates at the other endpoint $\big(\tfrac{1}{3-2t},1\big)$. First, we observe the following unboundedness result.
@@
On the other hand, by [31, Theorem 1.12], one has
@@
First, recall that for any $z,w\in{\mathbb{D}}$, there exists a Carleson tent $Q_{J}$, containing both $z$ and $w$ such that $|Q_{J}|\simeq\left|1-z\overline{w}\right|^{2}$ (see, e.g., [25, Section 3]). Then using the $1/3$-trick, one can find some $I\in{\mathcal{D}}$ or $\widetilde{{\mathcal{D}}}$ (see, Remark 1.4), such that
@@ 5. Hypersingular sparse operators
    $E(Q)\cap Q^{\prime}=\emptyset$ for any $^{\prime}\subset Q, QQ$$^{\prime}\in Q$.
Remark 5.2.
We remark that the above notion of sparse families was introduced by Lerner, Lorist, and Ombrosi [18] in their work on operator-free sparse domination. Compared with the usual definition used in the literature, this formulation is slightly more restrictive, most through the third condition above, which we refer to as the contracting property. A simple observation shows that this contracting property forces the sets $\{E(Q)\}_{Q\in\mathcal{S}}$ to be pairwise disjoint, thereby recovering the usual disjointness requirement in the standard definition of a sparse family.
Such sparse collections arise naturally in many applications of sparse domination, for instance: (1) collections of Carleson boxes; and (2) sparse collections produced by stopping-time constructions, such as those associated with maximal operators and Calderón–Zygmund operators. We refer the reader to [18] for further discussion and applications of sparse families with the contracting property.
We begin with the following model case. Let $t>1$, and for simplicity, assume that $\mathcal{S}$ be an $\eta$-sparse family of cubes in $\mathbb{R}^{n}$ such that $Q\subseteq Q_{0}=[0,1]^{n}$ for all $Q\in{\mathcal{S}}$. We define the (pointwise) hypersingular sparse operator associated with $\mathcal{S}$ by
@@ 5.1. Graded family
$\bullet$ First, let ${\mathcal{G}}^{(0)}$ denote the collection of all maximal dyadic cubes in $\mathcal{G}$}. For simplicity, we may assume that ${\mathcal{G}}^{(0)}=\{[0,1]^{n}\}$. Otherwise, we decompose $\mathcal{G}$} into finitely many such collections and treat each one separately (by translation and dilation). To this end, we define $\mathfrak{G}_{0}:=\ell(Q_{0})$, where $\ell(Q)$ denotes the sidelength of a dyadic cube $Q$.
$\bullet$ Next, let ${\mathcal{G}}^{(1)}$ be the subcollection of all maximal dyadic cubes in ${\mathcal{G}}\setminus{\mathcal{G}}^{(0)}$, and define $\mathfrak{G}_{1}:=\inf_{Q\in{\mathcal{G}}^{(1)}}\ell(Q)$. Iterating this procedure, we obtain a decomposition of $\mathcal{G}$} into layers $\{{\mathcal{G}}^{(j)}\}_{j\geq 0}$ together with the associated scales $\{\mathfrak{G}_{j}\}_{j\geq 0}$. Observe that for each $j\geq 1$, the dyadic cubes in ${\mathcal{G}}^{(j)}$ are mutually disjoint.
@@
Let ${\mathcal{G}}\subseteq{\mathcal{D}}$ be a collection of dyadic cubes in $\mathbb{R}^{n}$ such that ${\mathcal{G}}^{(0)}=\{[0,1]^{n}\}$, and write ${\mathcal{G}}=\bigcup_{j\geq 0}{\mathcal{G}}^{(j)}$ as above. We say that $\mathcal{G}$} is graded if
@@
Let $\eta\in(0,1)$, ${\mathcal{S}}\subseteq{\mathcal{D}}$ be a graded (contracting) $\eta$-sparse family in $0,1]^{n}$ with degree $K_{{\mathcal{S}}}$ as in Definition 5.5, $1<t<1-\frac{\log_{2}(1-\eta)}{nK_{{\mathcal{S}}}}$, and $\mathbb{A}_{{\mathcal{S}}}^{t}$ be the associated hypersingular sparse operator. Then the following statements hold.
@@
Observe that, by the sparseness assumption and contracting property of $\mathcal{S}$}, for any $j\geq 0$, one has
(5.4) $\sum_{Q\in{\mathcal{S}}^{(j)}}|Q|\lesssim(1-\eta)^{j},$
@@
(1). First, consider ${\mathcal{S}}^{(0)}=\{[0,1]^{n}\}$. Observe that
[…]
By the sparseness and contracting property of $\mathcal{S}$}, $\mathbb{A}_{{\mathcal{S}}}^{t}f(z)$ can only take larger values on $\bigcup_{Q\in{\mathcal{S}}^{(1)}}Q$, whose size is at most $1-\eta$.
(2). Next, we consider the next layer ${\mathcal{S}}^{(1)}$. Using the sparseness and contracting property of $\mathcal{S}$} again, we find that
$\mathbb{A}_{{\mathcal{S}}}^{t}f(z)\leq 1+\frac{1}{|Q|^{t}}\int_{Q}|f(x)|dx\leq 1+|Q|^{1-t}\leq 1+2^{nK_{{\mathcal{S}}}(t-1)},\quad z\in[0,1]^{n}\backslash\left(\bigcup_{Q\in{\mathcal{S}}^{(2)}}Q\right)$
and $\mathbb{A}_{{\mathcal{S}}}^{t}f(z)$ can only take larger values on $\bigcup_{Q\in{\mathcal{S}}^{(2)}}Q$, whose size is at most $(1-\eta)^{2}$.
@@
where $C_{1},C_{2}>0$ are some absolute constants [^6] Here, we can take $C_{1}=\frac{2^{nK_{{\mathcal{S}}}(t-1)}}{2^{nK_{{\mathcal{S}}}(t-1)}-1}$ and $C_{2}=\frac{1}{2^{nK_{{\mathcal{S}}}(t-1)}-1}$. that only depend on $n,K_{{\mathcal{S}}}$ and $t$, then $z\in[0,1]^{n}\backslash\left(\bigcup_{Q\in{\mathcal{S}}^{(J+1)}}Q\right)$, whose size is at least $1-(1-\eta)^{J+1}$. Therefore, for any $\alpha>0$ sufficiently large, if $\mathbb{A}_{{\mathcal{S}}}^{t}f(z)>C_{1}\alpha-C_{2}$, then
$z\in\bigcup_{Q\in{\mathcal{S}}^{(\widetilde{J}+1)}}Q$
@@ 6. Some open problems
where ${\mathbb{B}}_{n}$ is the unit ball in $\mathbb{C}^{n}$ and $dV$ is the normalized volume measure on ${\mathbb{B}}_{n}$. In [28], Zhao and Zhou characterized the strong-type $L^{p}_{\alpha}({\mathbb{B}}_{n})\to L^{q}_{\beta}({\mathbb{B}}_{n})$ bounds for these operators under various assumptions on the parameters $a,b,c,\alpha$, and $\beta$. Here, for $1\leq p<\infty$ and $1<\alpha<\infty$, the space $L^{p}_{\alpha}({\mathbb{B}}_{n})$ is defined with respect to the measure $dv_{\alpha}(z)=c_{\alpha}(1-|z|^{2})^{\alpha}\,dV(z)$, where $c_{\alpha}$ is chosen so that $v_{\alpha}({\mathbb{B}}_{n})=1$. A natural question is whether the method developed in the present paper can be applied to obtain critical-line estimates for Forelli–Rudin type operators in the hypersingular regime (see, e.g., [28, Theorem 1.1]).